# Realizable Circuit Complexity: Embedding Computation in Space–Time

### v2

@@ C Position in the History of Computational Theory
From information to realization. In 1948, Shannon’s information theory [20] established that the fundamental limits of communication are governed not by hardware, but by probability and entropy. Shannon’s framework abstracted away geometry and energy, defining an informational capacity that could, in principle, be realized by any physical medium.
Two decades later, Landauer and Bennett [21] reintroduced physics to information: computation, they argued, must obey thermodynamic laws. Landauer’s principle—that each logically irreversible operation dissipates at least $k_{B}T\ln 2$ joules of heat—marked the beginning of physical computability. This work recognized that information is not merely symbolic but embodied, subject to conservation and dissipation in the same way as energy.
@@
D Future Extensions to Other Domains
The framework of $\mathbf{RC}_{d}$ is not confined to Boolean or quantum circuits; it provides a general account of physically realizable information processing. Because its axioms arise from geometry (finite density), causality (bounded propagation), and thermodynamics (finite entropy flux), the same bounds apply to any substrate that computes through physical interaction and state evolution. Below we highlight how these constraints manifest across diverse computational domains.
Neuromorphic, embodied, and mortal computation.
Neural and embodied systems compute through continuous, spatially extended dynamics constrained by finite conduction speed and local connectivity. Within $\mathbf{RC}_{3}$, the flux limit $w(t)=\mathcal{O}(t^{2})$ imposes a ceiling on global synchronization and long-range coordination, forcing large-scale inference to emerge hierarchically. This constraint renders cortical and neuromorphic architectures—notably predictive-coding [22, 23] and columnar models [24, 25, 26, 27]—realizable in the physical sense: they operate near the entropy–flux bound of $\mathbf{RC}_{3}$. From this perspective, embodied intelligence [28] represents computation that is spatially embedded within causal geometry, while mortal computation [29] denotes the same process under finite energy and entropy budgets. Both are natural instantiations of realizable computation, grounded in the same physical limits of signal propagation and entropy flux.
Quantum and reversible computation.
Quantum systems obey the same geometric constraints while preserving information reversibly. The inclusion $\mathbf{RC}_{d}\subseteq\mathbf{QRC}_{d}$ follows from the existence of unitary embeddings for any classical process, but the converse fails because irreversible logic cannot in general be uncomputed. Even quantum systems are flux-bounded by the Lieb–Robinson limit [30, 31], which restricts information propagation to a linear light-cone. Thus $\mathbf{QRC}_{d}$ refines $\mathbf{BQP}$ by adding explicit spatial and flux constraints, clarifying why even quantum hardware cannot surpass the $t^{d-1}$ throughput limit [32, 33].
VLSI and spatial hardware design.
For planar devices ($d=2$), the $\mathbf{RC}_{2}$ bound recovers the area–time tradeoff $AT^{2}=\Omega(N^{2})$ of Thompson’s VLSI theory [6] and Rent’s empirical rule on interconnect scaling [7, 11]. Here $A=\mathcal{O}(t^{2})$ corresponds to physical area, while the boundary throughput $w(t)=\mathcal{O}(t)$ constrains wiring density. Hence $\mathbf{RC}_{d}$ subsumes classical VLSI theory as its two-dimensional limit and naturally extends it to higher-dimensional substrates [5].
Distributed, edge, and communication systems.
In distributed computing or sensor networks embedded in $\mathbb{R}^{d}$, information exchange between nodes is limited by the same causal and flux constraints. The $\mathcal{O}(t^{d-1})$ bound on cross-boundary transfer defines a geometric analogue of Shannon capacity [20], setting a hard limit on aggregate throughput. This result parallels Gupta and Kumar’s spatial capacity law for wireless networks, where total achievable rate scales as $O(\sqrt{n})$ in planar space [34]. Thus $\mathbf{RC}_{d}$ unifies circuit and network theories through a common geometric formulation of communication.
Energy-aware and thermodynamic machine learning.
When learning dynamics are interpreted as physical state updates, Liouville’s theorem and Gibbs entropy conservation imply that the energy–delay product of any model is lower-bounded by its entropy flux. Consequently, the description-length or free-energy minimization objectives in modern energy-based and predictive-coding models [35, 36, 37, 38, 23] are realizable precisely when their updates satisfy the $\mathbf{RC}_{d}$ bounds. This connects the Landauer limit [39, 40], information geometry, and computational generalization under a single physical principle.
Unified scaling law.
$t=\Omega\big(n^{1/(d-1)}\big),\qquad w(t)=\mathcal{O}(t^{d-1}),$

### v3

@@ C Final Remarks
From information to realization. In 1948, Shannon’s information theory [shannon1948mathematical] established that the fundamental limits of communication are governed not by hardware, but by probability and entropy. Shannon’s framework abstracted away geometry and energy, defining an informational capacity that could, in principle, be realized by any physical medium.
Two decades later, Landauer and Bennett [bennett1985fundamental] reintroduced physics to information: computation, they argued, must obey thermodynamic laws. Landauer’s principle—that each logically irreversible operation dissipates at least $k_{B}T\ln 2$ joules of heat—marked the beginning of physical computability. This work recognized that information is not merely symbolic but embodied, subject to conservation and dissipation in the same way as energy.